# Enhancing Delta Compression in LLMs via SVD-based Quantization Error Minimization

### v2

@@ 4.3 Analysis of $f_{\mathrm{max}}$
In DeltaMix, we set a hyperparameter termed $f_{\mathrm{max}}$ to constrain the number of active bit-widths during quantization. This section examines the performance of DeltaMix under varying values of $f_{\mathrm{max}}$. As shown in Table 3, DeltaMix consistently achieves better performance than Delta-CoMe across all settings, indicating that DeltaMix is insensitive to the choice of $f_{\mathrm{max}}$. In the main experiment, we set $f_{\mathrm{max}}$ to 4 to be consistent with Delta-CoMe.
@@ 4.4 Ablation of RTC
DeltaMix 62.7(0.04) 72.1(0.18) 80.2(2.09) 31.1(3.81) 61.5
DeltaMix (W/O RTC) 62.8(0.02) 72.2(0.05) 78.2(0.28) 27.5(3.81) 60.2
We conducted experiments to assess the necessity of RTC, as detailed in Table 4. Overall, RTC consistently enhances our method, yielding an average performance improvement of 2.2%. The results indicate that mitigating the deviation in the quantization loss of $\mathbf{U}$ enables DeltaMix to retain more information from $\Delta\mathbf{W}$. The importance of RTC is particularly pronounced in challenging tasks; for instance, it improves performance by 13.1% on the AIME2024 task. This improvement can be attributed to the more substantial quantization errors associated with quantizing $\mathbf{V}$ in these cases, thereby highlighting the critical need for reconstruction target correction.
@@ 5 Analyses
5.1 Inference Speed and Memory Cost
As shown in Figure 3 (left), a single GPU can deploy only two aligned models simultaneously. In contrast, it can support up to 8 and 12 models concurrently for Delta-CoMe and DeltaMix, respectively. This enhancement is attributable to the fact that, as the number of models increases, both methods necessitate only the additional deployment of compressed delta weights, thereby significantly reducing memory overhead. Notably, while Delta-CoMe exhausts GPU memory at 12 models, DeltaMix does not. Our further analysis indicates that DeltaMix typically employs fewer ranks, namely allocates a greater number of singular vectors with a bid-width of 0, thereby enhancing the GPU memory utilization efficiency.
For the end-to-end decoding latency illustrated in Figure 3 (middle, right), we find that Delta-CoMe and DeltaMix introduce overhead to Naive when the number of deployed models is small. However, Delta-CoMe and DeltaMix scale better and effectively translate the saved GPU memory into improved decoding latency. In contrast, the Naive approach quickly encounters out-of-memory issues. Furthermore, DeltaMix exhibits a superior generation speed compared to Delta-CoMe at scale, while the prefill times for both methods remain comparable. In Appendix E.2, we conduct more latency evaluation under varying arrival rates and request distributions following (Yao et al., 2024).
@@ 5.1 Inference Speed and Memory Cost
In this section, we train the DeepSeek-LLM-7B-Base (DeepSeek-AI, 2024) on math and code tasks
Table 5: Performance comparison between Delta-Compression and LoRA. Aligned is full fine-tuned model. For DeltaMix, we report the results in the format “mean(std)” with three runs.
@@
DeltaMix 1/16 43.3(0.6) 50.2(0.82) 13.5(0.76) 56.1(0.82) 40.8
using both LoRA and full fine-tuning. We subsequently apply DeltaMix to the delta weights of the fully fine-tuned model. Additional experimental details can be found in Appendix D.2. Table 5 presents a comparison of DeltaMix with LoRA. The results indicate that DeltaMix consistently outperforms LoRA across all tasks. Notably, DeltaMix achieves an average score of 40.8, which is close to the aligned model’s score of 42.0, representing a 14.9% improvement over LoRA.
@@ 5.3 Analyzing Quantization Error
DeltaMix 0.66 1.46 0.66 1.12 6.81 426.20
To better understand the difference between various delta-compression methods, we compute the quantization error on Qwen2.5-Math-7B-Instruct model as defined in Equation (1). Since outliers play a critical role in model compression (Dettmers et al., 2023; Lin et al., 2024), we also report the average error for the top 1% of activations with the largest absolute values in the aligned model, categorizing them as outliers. As different layers contribute differently to the final output  (Wu et al., 2024), we categorize the first 9 layers, layers 9 to 17, and the last 10 layers as low, mid, and high groups, respectively, and report the average error of each group. See Table 11 of Appendix E.5 for more details.
As demonstrated in Table 6, DeltaMix consistently exhibits lower overall quantization error compared to all baseline methods, attributable to its inherent objective of minimizing quantization error. In the mid layers, DeltaMix shows a slightly higher error than BitDelta, with values of 0.66 versus 0.61 for all activations and 1.12 versus 1.08 for outlier activations, respectively. However, it is important to note that since BitDelta is an empirical method, it cannot guarantee low quantization error across all layers. For example, in the high layers, BitDelta exhibits significantly higher error rates compared to DeltaMix, with values of 21.51 versus 6.81 for all activations and 3162.58 versus 426.20 for outlier activations, respectively. These experiments further illustrate that DeltaMix effectively reduces quantization error, thereby preserving the information contained in the delta weights as much as possible. In Appendix E.4, we visualize the bit allocation results of DeltaMix across different weight types and layers using the Qwen2.5-Math-7B-Instruct model.
@@ 6 Conclusion
In this study, we present DeltaMix, an adaptive mixed-precision delta-compression framework aimed at minimizing quantization error in the SVD space without introducing additional assumptions. DeltaMix offers a theoretical proof of the necessity for mixed-precision delta-compression and provides a practical quantization solution that involves solving a 0/1 linear integer programming problem and employing a reconstruction target correction method. DeltaMix outperforms all baseline delta-compression methods across four distinct downstream tasks, including reasoning, math, code, and multi-modal tasks, utilizing eight widely adopted aligned LLMs with backbone pre-trained models, including Qwen2.5, Qwen2.5-Math, Qwen2.5-Coder, and LLaMA2. Moreover, DeltaMix significantly reduces deployment costs by minimizing memory overhead and accelerating inference. We believe that DeltaMix provides considerable theoretical and practical value, particularly in scenarios involving multi-tenant deployments.
Ethics statement
We propose an adaptive mixed-precision delta-compression framework designed to minimize quantization error in the singular value decomposition space. Our experiments rely exclusively on publicly available datasets and models, without involving human subjects or sensitive data. We do not anticipate any direct negative consequences arising from this approach.
Reproducibility statement
To facilitate reproducibility, we describe our experimental setup in Section 4.1 and provide additional details, including models, datasets, metrics, and GPUs, in Appendix D.
Appendix A Limitation and Broader Impact
DeltaMix significantly reduces hardware requirements and computational costs for serving multiple finetuned models, thereby enabling smaller entities to deploy advanced large language models more feasibly. Additionally, it lowers power consumption and reduces the carbon emissions associated with LLM deployment. Despite DeltaMix ’s demonstrated improvements over baseline methods in reducing the performance gap between compressed and aligned models, it is important to note that DeltaMix remains a lossy compression method for certain tasks. We believe this is an important consequence and encourage future research to further minimize this performance gap, particularly in tasks where performance degradation is substantial.
Appendix B LLMs Usage
In this work, large language models (LLMs) were used solely as auxiliary tools for grammar correction and text refinement.
Appendix C Formula Derivation
C.1 $\mathbf{V}$ Hessian Matrix
@@
C.2 $\mathbf{U}$ Hessian Matrix
@@
C.3 Detailed Derivation Process for new U
@@
By setting the gradient of the loss to zero, DeltaMix gets the corrected $\mathbf{\tilde{U}}$ as follow:
@@
Appendix D Experiments Setup
D.1 Main Experiments
Task 7B Models 13-14B Models
Backbone Aligned Backbone Aligned
Math Qwen2.5-Math Qwen2.5-Math-Instruct LLaMA2 MetaMath
Multi-Modal Qwen2.5 Qwen2.5-VL-Instruct LLaMA2 LLAVA-V1.5
@@ D.1 Main Experiments
$\bullet$ Reasoning. We use the Math500 and AIME2024 datasets as the test set. For the reasoning prompt of AIME2024, we follow with (Naman Jain & et al., 2024). The maximum length of both tasks is set to 8192. The evaluation metric is accuracy, determined by comparing the model-generated solution to the ground truth.
@@
$\bullet$ Multi-Modal. We utilize the GQA (Hudson & Manning, 2019) and the image part of ScienceQA (Lu et al., 2022) datasets, both commonly used for evaluating VLM performance, as our test set. We adopt lmms-eval (Zhang et al., 2024) to evaluate both tasks. The evaluation metric is accuracy, which measures whether the model selects the correct option.
To accelerate DeltaMix’s quantization, we discard the last $k$ ranks of $\mathbf{V}$, where $k=\left\lfloor\frac{G_{b}(h_{\mathrm{in}}\cdot h_{\mathrm{out}})}{(h_{\mathrm{in}}+h_{\mathrm{out}})\cdot\mathrm{bit_{min}}}\right\rfloor$. Here, $\mathrm{bit_{min}}$ denotes the smallest non-zero bit-width allowed in quantization. This strategy is motivated by the observation that larger singular values correspond to more important singular vectors. Due to storage constraints, the last $k$ singular vectors would be assigned 0-bit and thus excluded from optimization.
D.2 Delta-Compression vs. Delta-Tuning
Specifically, we set the LoRA rank to 128 and the scale factor to 128, training LoRA for all model parameters for 3 epochs using a cosine schedule with a peak learning rate of 4e-5 and a warm-up ratio of 0.1, using model deepseek-llm-7b-base (DeepSeek-AI, 2024). We randomly sample 50k training examples from MetaMathQA (Yu et al., 2023) and Magicoder-Evol-Instruct (Wei et al., 2024) for the math and code tasks, respectively. To ensure a fair comparison, we fine-tune all model parameters using the same datasets as those used for LoRA training. We then apply DeltaMix to both math and code finetuned LLMs.
Appendix E More Experiments
E.1 Analyzing the Different Quantization Schemes in U
Table 8: We evaluate the performance of various quantization schemes applied to $\mathbf{U}$ on Qwen2.5-Math-7B-Instruct. Here, “x-bit” denotes quantization of $\mathbf{U}$ at x-bit precision. The “DeltaMix-row” setting refers to applying the optimization model to determine the scheme and performing quantization in a row-wise manner, whereas “DeltaMix ” indicates employing the same quantization scheme used for $\mathbf{V}$, with quantization carried out column by column.
@@ E.1 Analyzing the Different Quantization Schemes in U
U(2bit),V(DeltaMix) 1/16 76.8 93.6 85.2
U(3bit),V(DeltaMix) 1/16 75.6 93.4 84.5
U(DeltaMix-row),V(DeltaMix) 1/16 75.2 93.6 84.4
DeltaMix 1/16 75.2 93.9 84.6
@@
In this section, we investigate the effect of applying different quantization schemes to $\mathbf{U}$ in order to assess the necessity of mixed precision. Our evaluation is conducted on Qwen2.5-Math-7B-Instruct. The results show that there is no significant difference between DeltaMix and other quantization methods for $\mathbf{U}$. As shown in Table 8, “x-bit” denotes quantization of $\mathbf{U}$ with x-bit precision. The “DeltaMix-row” setting applies the optimization model to determine the scheme and performs quantization in a row-wise manner, whereas “DeltaMix” adopts the same quantization scheme as $\mathbf{V}$ and conducts quantization column by column. The performance differences across schemes are minimal, with the largest gap in average scores being only 0.95%, observed between the “DeltaMix-row” setting and the 2-bit quantization. These results suggest that the choice of quantization strategy for $\mathbf{U}$ has only a limited impact on overall performance.
E.2 Inference Speed and Memory Cost

### v3

@@ 4.2 Main Results
When analyzing the various tasks, we observe that PrinMix exhibits more pronounced improvements in challenging scenarios characterized by a significant performance gap between the baseline methods and the aligned model. This is particularly evident in reasoning-intensive benchmarks, such as AIME2024, as well as in multimodal tasks utilizing 7B backbones. For instance, PrinMix surpasses the previous state-of-the-art model, Delta-CoMe, by 22.3% on the 7B model and by 26.9% on the 14B model. Further analysis reveals that these models display larger norms for $\Delta\mathbf{W}$. Specifically, the median norm of DeepSeek-R1-Distill-Qwen-7B and Qwen2.5-VL-Instruct is 6.5 and 10.3 times that of Qwen-Coder-Instruct-7B, with corresponding values of 26.13 and 41.45 compared to 4.02, respectively. In this context, baseline methods struggle to achieve optimal solutions due to their empirical nature. In contrast, PrinMix directly optimizes quantization error from a mathematical perspective, enabling it to fully leverage its strengths in demanding tasks. However, on tasks where baselines already achieve near-lossless accuracy, such as MBPP and HumanEval on the 7B backbone, PrinMix performs comparably to the best baseline. In these scenarios, the norm of $\Delta\mathbf{W}$ is relatively small and can be easily compressed, leading to a ceiling effect: $\Delta\mathbf{W}$ can be quantized almost losslessly by existing baselines, leaving little room for further improvement.
We also report the quantization time cost of PrinMix. Please refer to Appendix C.4 for more details. The results show that PrinMix requires only 1.1 hours for 7B models and 2.4 hours for 14B models on a single GPU. This time can be at least reduced by $2\times$ by switching to a commercial solver and reducing the solution space, as discussed in C.4.
4.3 Ablation of $f_{\mathrm{max}}$
In PrinMix, we set a hyperparameter termed $f_{\mathrm{max}}$ to constrain the number of active bit-widths during quantization. This section examines the performance of PrinMix under varying values of $f_{\mathrm{max}}$. As shown in Table 4, PrinMix consistently achieves better performance than Delta-CoMe across all settings, indicating that PrinMix is insensitive to the choice of $f_{\mathrm{max}}$. In the main experiment, we set $f_{\mathrm{max}}$ to 4 to be consistent with Delta-CoMe.
@@ 4.4 Ablation of RTC
PrinMix 62.7(0.04) 72.1(0.18) 80.2(2.09) 31.1(3.81) 61.5
PrinMix (W/O RTC) 62.8(0.02) 72.2(0.05) 78.2(0.28) 27.5(3.81) 60.2
We conducted experiments to assess the necessity of RTC, as detailed in Table 5. Overall, RTC consistently enhances our method, yielding an average performance improvement of 2.2%. The results indicate that mitigating the deviation in the quantization loss of $\mathbf{U}$ enables PrinMix to retain more information from $\Delta\mathbf{W}$. The importance of RTC is particularly pronounced in challenging tasks; for instance, it improves performance by 13.1% on the AIME2024 task. This improvement can be attributed to the more substantial quantization errors associated with quantizing $\mathbf{V}$ in these cases, thereby highlighting the critical need for reconstruction target correction.
@@ 5 Analyses
In this section, we train the DeepSeek-LLM-7B-Base (DeepSeek-AI, 2024) on math and code tasks using both LoRA and full fine-tuning. We subsequently apply PrinMix to the delta weights of the fully fine-tuned model and LoRA. Additional experimental details can be found in Appendix B.2. Table 6 presents a comparison of PrinMix with LoRA. PrinMix consistently outperforms LoRA across all tasks. When $\alpha=1/16$, PrinMix achieves an average score of 40.8, which is close to the aligned model’s score of 42.0, representing a 14.9% improvement over LoRA.
In addition to full fine-tuning model, PrinMix can effectively compress the LoRA model as well. As illustrated in Table 6, PrinMix-LoRA demonstrates a performance degradation of only 0.3 compared to the LoRA, while further compressing LoRA with a ratio of 4 ($\alpha=1/64$). Notably, Baselines like BitDelta and Delta-CoMe cannot apply to LoRA. BitDelta directly quantizes $\Delta\mathbf{W}$ to 1 bit without employing any low-rank approximation. Consequently, it cannot effectively utilize the low-rank properties inherent in LoRA. For Delta-CoMe, the empirically determined mixed-precision scheme is fixed and does not offer a clear method for allocating mixed precision at other compression ratios. In contrast, PrinMix allows compression of $\Delta\mathbf{W}$ to arbitrary ratios, making it more flexible and practically advantageous. Although PrinMix-LoRA performs well on top of LoRA, it underperforms PrinMix at the same compression ratio ($\alpha=1/64$). This further highlights the significance of delta-compression in comparison to delta-tuning.
Table 6: Performance comparison between Delta-Compression and LoRA. Aligned is full fine-tuned model. For PrinMix, we report the results in the format “mean(std)” with three runs.
@@
PrinMix 1/16 43.3(0.60) 50.2(0.82) 13.5(0.76) 56.1(0.82) 40.8
PrinMix-LoRA 1/64 34.1(1.64) 47.7(0.85) 9.1(0.50) 49.7(0.30) 35.2
PrinMix 1/64 39.4(1.56) 50.0(1.10) 11.1(0.95) 52.9(0.57) 38.4
5.2 Inference Speed and Memory Cost
As shown in Figure 3 (left), a single GPU can deploy only two aligned models simultaneously. In contrast, it can support up to 8 and 12 models concurrently for Delta-CoMe and PrinMix, respectively. This enhancement is attributable to the fact that, as the number of models increases, both methods necessitate only the additional deployment of compressed delta weights, thereby significantly reducing memory overhead. Notably, while Delta-CoMe exhausts GPU memory at 12 models, PrinMix does not. Our further analysis indicates that PrinMix typically employs fewer ranks, namely allocates a greater number of singular vectors with a bid-width of 0, thereby enhancing the GPU memory utilization efficiency.
For the end-to-end decoding latency illustrated in Figure 3 (middle, right), we find that Delta-CoMe and PrinMix introduce overhead to Naive when the number of deployed model is small. However, Delta-CoMe and PrinMix scale better and effectively translate the saved GPU memory into improved decoding latency. In contrast, the Naive approach quickly encounters out-of-memory issues. Furthermore, PrinMix exhibits a superior generation speed compared to Delta-CoMe at scale, while the prefill times for both methods remain comparable. In Appendix C.6, we conduct more latency evaluation under varying arrival rates and request distributions following DeltaZip (Yao et al., 2024).
@@ 5.3 Analyzing Quantization Error
PrinMix 0.66 1.46 0.66 1.12 6.81 426.20
To better understand the difference between various delta-compression methods, we compute the quantization error on Qwen2.5-Math-7B-Instruct model as defined in Equation (1). Since outliers play a critical role in model compression (Dettmers et al., 2023; Lin et al., 2024), we also report the average error for the top 1% of activations with the largest absolute values in the aligned model, categorizing them as outliers. As different layers contribute differently to the final output  (Wu et al., 2024), we categorize the first 9 layers, layers 9 to 17, and the last 10 layers as low, mid, and high groups, respectively, and report the average error of each group. See Table 15 of Appendix C.8 for more details.
As demonstrated in Table 7, PrinMix consistently exhibits lower overall quantization error compared to all baseline methods, attributable to its inherent objective of minimizing the quantization error. In the mid layers, PrinMix shows a slightly higher error than the BitDelta, with values of 0.66 versus 0.61 for all activations and 1.12 versus 1.08 for outlier activations, respectively. However, it is important to note that since BitDelta is an empirical method, it cannot guarantee low quantization error across all layers. For example, in the high layers, BitDelta exhibits significantly higher error rates compared to PrinMix, with values of 21.51 versus 6.81 for all activations and 3162.58 versus 426.20 for outlier activations, respectively. These experiments further illustrate that PrinMix effectively reduces quantization error, thereby preserving the information contained in the delta weights as much as possible. In Appendix C.7, we visualize the bit allocation results of PrinMix across different weight types and layers using the Qwen2.5-Math-7B-Instruct model.
@@ 6 Conclusion
In this study, we present PrinMix, an adaptive mixed-precision delta-compression framework aimed at minimizing quantization error in the SVD space without introducing additional assumptions. PrinMix offers a theoretical proof of the necessity for mixed-precision delta-compression and provides a practical quantization solution that involves solving a 0/1 linear integer programming problem and employing a reconstruction target correction method. PrinMix outperforms all baseline delta-compression methods across four distinct downstream tasks, including reasoning, math, code, and multi-modal, utilizing eight widely adopted aligned LLMs with backbone pre-trained models, including Qwen2.5, Qwen2.5-Math, Qwen2.5-Coder, and LLaMA2. Moreover, PrinMix significantly reduces deployment costs by minimizing memory overhead and accelerating inference. We believe that PrinMix provides considerable theoretical and practical value, particularly in scenarios involving multi-tenant deployments.
Impact Statement
PrinMix significantly reduces hardware requirements and computational costs for serving multiple finetuned models, thereby enabling smaller entities to deploy advanced large language models more feasibly. Additionally, it lowers power consumption and reduces the carbon emissions associated with LLM deployment. Despite PrinMix ’s demonstrated improvements over baseline methods in reducing the performance gap between compressed and aligned models, it is important to note that PrinMix remains a lossy compression method for certain tasks. We believe this is an important consequence and encourage future research to further minimize this performance gap, particularly in tasks where performance degradation is substantial.
Appendix A Formula Derivation
A.1 $\mathbf{V}$ Hessian Matrix
@@
A.2 $\mathbf{U}$ Hessian Matrix
@@
A.3 Detailed Derivation Process for new U
@@
By setting the gradient of the loss to zero, PrinMix gets the corrected $\mathbf{\tilde{U}}$ as follow:
@@
Appendix B Experiments Setup
B.1 Main Experiments
@@
$\bullet$ Reasoning. We use the Math500 and AIME2024 datasets as the test set. For the reasoning prompt of AIME2024, we follow with (Naman Jain and et al., 2024). The maximum length of both tasks is set to 8192. The evaluation metric is accuracy, determined by comparing the model-generated solution to the ground truth.
@@
$\bullet$ Multi-Modal. We utilize the GQA (Hudson and Manning, 2019) and the image part of ScienceQA (Lu et al., 2022) datasets, both commonly used for evaluating VLM performance, as our test set. We adopt lmms-eval (Zhang et al., 2024) to evaluate both tasks. The evaluation metric is accuracy, which measures whether the model selects the correct option.
B.2 Delta-Compression vs. Delta-Tuning
Specifically, we set the LoRA rank to 128 and the scale factor to 256, training LoRA for all model parameters for 3 epochs using a cosine schedule with a peak learning rate of 4e-5 and a warm-up ratio of 0.1, using model deepseek-llm-7b-base (DeepSeek-AI, 2024). We randomly sample 50k training examples from MetaMathQA (Yu et al., 2023) and Magicoder-Evol-Instruct (Wei et al., 2024) for the math and code tasks, respectively. To ensure a fair comparison, we fine-tune all model parameters using the same datasets as those used for LoRA training. We then apply PrinMix to both math and code finetuned LLMs.
Appendix C More Experiments
C.1 Analyzing the Different Quantization Schemes in U
Table 8: We evaluate the performance of various quantization schemes applied to $\mathbf{U}$ on Qwen2.5-Math-7B-Instruct. Here, “x-bit” denotes quantization of $\mathbf{U}$ at x-bit precision. The “PrinMix-row” setting refers to applying the optimization model to determine the scheme and performing quantization in a row-wise manner, whereas “PrinMix ” indicates employing the same quantization scheme used for $\mathbf{V}$, with quantization carried out column by column.
@@ C.1 Analyzing the Different Quantization Schemes in U
U(2bit),V(PrinMix) 1/16 76.8 93.6 85.2
U(3bit),V(PrinMix) 1/16 75.6 93.4 84.5
U(PrinMix-row),V(PrinMix) 1/16 75.2 93.6 84.4
PrinMix 1/16 75.2 93.9 84.6
@@
In this section, we investigate the effect of applying different quantization schemes to $\mathbf{U}$ in order to assess the necessity of mixed precision. Our evaluation is conducted on Qwen2.5-Math-7B-Instruct. The results show that there is no significant difference between PrinMix and other quantization methods for $\mathbf{U}$. As shown in Table 8, “x-bit” denotes quantization of $\mathbf{U}$ with x-bit precision. The “PrinMix-row” setting applies the optimization model to determine the scheme and performs quantization in a row-wise manner, whereas “PrinMix” adopts the same quantization scheme as $\mathbf{V}$ and conducts quantization column by column. The performance differences across schemes are minimal, with the largest gap in average scores being only 0.95%, observed between the “PrinMix-row” setting and the 2-bit quantization. These results suggest that the choice of quantization strategy for $\mathbf{U}$ has only a limited impact on overall performance.
C.2 Ablation of Compression Ratio
To show that PrinMix can apply to arbitrary compression ratios, we evaluate Qwen2.5-Math-7B-Instruct at four compression ratios, as shown in Table 9. The performance of PrinMix decreases as the compression ratio increases. This is expected, as
Table 9: Performance of PrinMix under different compression ratios $1/\alpha$.
Compression Ratio DeepSeek-R1-Distill-Qwen-7B Qwen2.5-Math-7B-Instruct
$\alpha$ Math500 AIME2024 Math500 GSM8K AVG
3/16 86.4 36.7 77.2 95.6 74.0
2/16 85.8 33.3 77.4 95.1 72.9
1/16 83.2 33.3 77.6 94.8 72.2
1/32 76.8 26.7 73.4 91.6 67.1
a higher compression ratio indicates a reduced capacity of the quantized model to preserve information from the original model. Notably, baselines like BitDelta and Delta-CoMe cannot apply to other compression ratios except $\alpha=$1/16. BitDelta quantizes $\Delta\mathbf{W}$ to a fixed 1 bit, resulting in a constant compression ratio. For Delta-CoMe, the empirically determined mixed-precision scheme is fixed and does not offer a clear method for allocating mixed precision at other compression ratios. In contrast, PrinMix enables the compression of $\Delta\mathbf{W}$ to arbitrary ratios, offering greater flexibility and broader applicability.
Table 10: The performance of PrinMix to quantize Qwen2.5-Math-7B-Instruct with different number of calibration data.
Calibration Size Math500 GSM8K AVG
16 76.4 94.5 85.5
32 76.2 95.1 85.7
64 76.8 94.3 85.6
128 77.6 94.8 86.2
256 76.0 94.1 85.1
Table 11: The performance of PrinMix to quantize Qwen2.5-Math-7B-Instruct using calibration data drawn from C4 and Wikitext2.
Math500 GSM8K AVG
C4 77.6 94.8 86.2
Wikitext2 76.6 94.8 85.7
MetaMath 75.4 93.6 84.5
C.3 Ablation of Calibration Dataset
Since PrinMix is a calibration-dependent method, to verify its robustness on calibration, we conducted experiments with different sizes and domains of the calibration dataset to quantize Qwen2.5-Math-7B-Instruct. For calibration on domains, each calibration set contains 128 randomly sampled sequences of length 2048. Due to the insufficient number of sequences of this length in the MetaMathQA dataset, we concatenated multiple question–answer pairs in a few-shot format. To examine the effect of dataset size on calibration, we varied the number of calibration samples on C4 dataset from 16 to 256. The results in Tables 11 and 11 demonstrate that PrinMix performs well on all calibration setups. The average performance gap is within 2.0%, confirming PrinMix ’s robustness.
C.4 Time For Quantization
Simulation Optimization Quantization Total
PrinMix Q_proj 3.1 17.5 1.0 134.3
K_proj 3.1 1.0
V_proj 3.1 1.0
O_proj 4.0 11.5 1.5
Up_proj 3.3 20.5 2.8
Gate_proj 3.3 2.8
Down_proj 19.7 24.1 11.0
In this section, we evaluate the quantization time of PrinMix within a single transformer block. PrinMix determines the mixed-precision quantization strategy by minimizing quantization loss, formulated as a 0/1 integer linear programming problem. To clarify the computational overhead, we decompose the quantization time into three components. The first is “simulation time”, which reflects the cost of estimating quantization loss under different bit-widths. The second is “optimization time”, incurred when solving the 0/1 integer linear programming problem. The third is the “quantization time” itself, representing the cost of quantizing each linear layer according to the selected strategy. The corresponding results for one transformer block of Qwen2.5-Math-7B-Instruct, which contains 28 blocks in total, are summarized in Table 12. As shown in Table  12, most of the quantization time is spent on “simulation” and “optimization”. For example, the “simulation” and “optimization” times in Q, K, and V_proj are 9.3s and 17.5s, respectively, while the quantization time only takes 3.0s. Specifically, simulation time increases with the number of columns, while optimization time grows with the number of rows.
Overall, although PrinMix requires 1.1 hours for 7B models and 2.4 hours for 14B models on a single L20 GPU, this is acceptable.
Accelerating the Optimization Time.
As illustrated in Table 12, solving the optimization problem constitutes the primary bottleneck, dominating the total quantization time. To mitigate this overhead, we propose accelerating the ILP solving process through two main avenues: ➀ increasing the ILP solving speed, and ➁ constraining the solution space. Specifically, regarding ➀, while we employ open-source solvers in this study, switching to a commercial solver like COPT (Ge et al., 2023) can yield a $6\times$ speedup. Furthermore, for ➁, the process can be accelerated by limiting the number of candidate bit-widths (e.g., reducing candidates from 8 to 4). Our experiments in Appendix C.5 demonstrate that PrinMix performs consistently well with different numbers of candidate bit-widths.
These methods significantly enhance PrinMix’s efficiency in practical deployment. Leveraging the acceleration strategies yields a 4x speedup in solving the optimization problem. Consequently, the quantization process for PrinMix only takes 0.5 hours for 7B models and 1.2 hours for 14B models, representing a 50% reduction in temporal overhead.
C.5 Sensitivity to the Number of Candidate Bit-widths
Table 13: Performance of PrinMix using incrementally expanded candidate bit-widths.
Candidate Bit-widths Qwen2.5-Math-7B-Instruct
Math500 GSM8K AVG
$\{0,2,3,4\}$ 76.7 94.2 85.5
$\{0,2,3,4,5\}$ 77.1 94.0 85.7
$\{0,2,3,4,5,6\}$ 76.9 94.5 85.5
$\{0,2,3,4,5,6,7\}$ 77.3 94.4 85.9
$\{0,2,3,4,5,6,7,8\}$ 77.7 94.6 86.2
In this section, we investigate the sensitivity of PrinMix in the number of candidate bit-widths. We construct the candidates incrementally from {0, 2, 3, 4} to {0, 2, 3, 4, 5, 6, 7, 8} and quantize Qwen2.5-Math-7B-Instruct at $\alpha=1/16$ in these 5 different configurations. As shown in Table 13, the performance gap between the best and worst average score is 0.6%, which demonstrate that PrinMix is robust on the number of candidate bit-widths.
C.6 Inference Speed and Memory Cost